# Layer-adaptive Expert Pruning for Pre-Training of Mixture-of-Experts Large Language Models

### v1

@@ Abstract
Although Mixture-of-Experts (MoE) Large Language Models (LLMs) deliver superior accuracy with a reduced number of active parameters, their pre-training represents a significant computationally bottleneck due to underutilized experts and limited training efficiency. This work introduces a Layer-Adaptive Expert Pruning (LAEP) algorithm designed for the pre-training stage of MoE LLMs. In contrast to previous expert pruning approaches that operate primarily in the post-training phase, the proposed algorithm enhances training efficiency by selectively pruning underutilized experts and reorganizing experts across computing devices according to token distribution statistics. Comprehensive experiments demonstrate that LAEP effectively reduces model size and substantially improves pre-training efficiency. In particular, when pre-training the 1010B Base model from scratch, LAEP achieves a 48.3% improvement in training efficiency alongside a 33.3% parameter reduction, while still delivering excellent performance across multiple domains.
@@ 1 Introduction
  • Apply LAEP to pre-train a 1515B-parameter sparse MoE model, achieving a 33.3% reduction in model parameters and a 48.3% improvement in training efficiency compared with the base model. Moreover, the resulting 1010B pre-trained base model attains performance comparable to state-of-the-art systems on major benchmarks across diverse domains.
@@ 3.4.4 Scalability of LAEP
4 Pre-training of 1010B Base model with LAEP
@@
Table 5: Performance comparison of the base model 1515B (activated 68.5B) under different methods. The tests are conducted on 824 AI chips, while the numerical precision during the training process is set to BF16. The coefficient for all auxiliary losses is set to 0.01, a value significantly higher than standard practice, to ensure near-ideal expert load balancing. Regarding the LAEP hyperparameters, $\alpha$ is set to 0.2 for the first and last one-sixth of the layers and 0.4 for the remaining layers, while $\beta$ is fixed at 0.1.
@@
Table 6: Comparison on tasks between the 1010B Base model and other representative dense and MoE base models.
Benchmark #Shots LLaMA-3.1-405B Base DeepSeek-V3-Base 1010B Base
@@
We compare the accuracy of 1010B Base model with that of the DeepSeek-V3-Base model and the Llama-3.1-405B Base model (Table 6). It can be observed that our model achieves leading accuracy on the Natural Questions, HumanEval and MBPP code evaluation tasks, as well as the Math mathematics task. On the GSM8K math task, the accuracy of our model is lower than that of DeepSeek-V3-Base but higher than that of Llama-3.1-405B. Additionally, on the Pile-test, MMLU, and ARC-Challenge, the accuracy of our model is lower than that of both DeepSeek-V3-Base and LLaMA-3.1-405B. Overall, the 1010B model pre-trained using the LAEP method achieves accuracy comparable to that of DeepSeek-V3-Base across various tasks, validating the effectiveness of our approach.
@@ A.2 Text Corpus for Pre-training
The training datasets for the 10B and 20B models consist of 100B tokens, with a test set of 10B tokens. The 1010B Base model is trained on 1080B tokens. All the aforementioned data are uniformly sampled from the pre-training datasets introduced in this section.

### v2

@@ Abstract
Although Mixture-of-Experts (MoE) Large Language Models (LLMs) deliver superior accuracy with a reduced number of active parameters, their pre-training represents a significant computationally bottleneck due to underutilized experts and limited training efficiency. This work introduces a Layer-Adaptive Expert Pruning (LAEP) algorithm designed for the pre-training stage of MoE LLMs. In contrast to previous expert pruning approaches that operate primarily in the post-training phase, the proposed algorithm enhances training efficiency by selectively pruning underutilized experts and reorganizing experts across computing devices according to token distribution statistics. Comprehensive experiments demonstrate that LAEP effectively reduces model size and substantially improves pre-training efficiency. In particular, when pre-training the Yuan3.0-1T Base model from scratch original with 1515B parameters, LAEP achieves a 48.3% improvement in training efficiency alongside a 33.3% parameter reduction, while still delivering excellent performance across multiple domains.
@@ 1 Introduction
  • Apply LAEP to pre-train a 1515B-parameter sparse MoE model, achieving a 33.3% reduction in model parameters and a 48.3% improvement in training efficiency compared with the base model. Moreover, the resulting Yuan3.0-1T pre-trained base model attains performance comparable to state-of-the-art systems on major benchmarks across diverse domains.
@@ 3.4.4 Scalability of LAEP
4 Pre-training of Yuan3.0-1T Base model with LAEP
@@
Table 5: Performance comparison of the Yuan3.0-1T Base model 1515B (activated 68.5B) under different methods. The tests are conducted on 824 AI chips, while the numerical precision during the training process is set to BF16. The coefficient for all auxiliary losses is set to 0.01, a value significantly higher than standard practice, to ensure near-ideal expert load balancing. Regarding the LAEP hyperparameters, $\alpha$ is set to 0.2 for the first and last one-sixth of the layers and 0.4 for the remaining layers, while $\beta$ is fixed at 0.1.
@@
Table 6: Comparison on tasks between the Yuan3.0-1T Base model and other representative dense and MoE base models.
Benchmark #Shots LLaMA-3.1-405B Base DeepSeek-V3-Base Yuan3.0-1T Base
@@
We compare the accuracy of Yuan3.0-1T Base model with that of the DeepSeek-V3-Base model and the Llama-3.1-405B Base model (Table 6). It can be observed that our model achieves leading accuracy on the Natural Questions, HumanEval and MBPP code evaluation tasks, as well as the Math mathematics task. On the GSM8K math task, the accuracy of our model is lower than that of DeepSeek-V3-Base but higher than that of Llama-3.1-405B. Additionally, on the Pile-test, MMLU, and ARC-Challenge, the accuracy of our model is lower than that of both DeepSeek-V3-Base and LLaMA-3.1-405B. Overall, the Yuan3.0-1T model pre-trained using the LAEP method achieves accuracy comparable to that of DeepSeek-V3-Base across various tasks, validating the effectiveness of our approach.
@@ A.2 Text Corpus for Pre-training
The training datasets for the 10B and 20B models consist of 100B tokens, with a test set of 10B tokens. The Yuan3.0-1T Base model is trained on 1080B tokens. All the aforementioned data are uniformly sampled from the pre-training datasets introduced in this section.